A positive charge of 4.80 μC ís fixed in place. From a distance of 4.30 cm a particle of mass 6.20 g and charge +3.70 μC is fired with an initial speed of 64.0 m/s directly toward the fixed charge. How close to the fixed charge does the particle get before it comes to rest and starts traveling away?
Answer #1

Using conservation of energy

Change in electric potential Energy of particle = change in KE of particle

k q1 q2 ( 1/ r1 - 1/r2 ) = 0.5* m (v^2 - 0)

9*10^9* 4.8*10^-6*3.7*10^-6* (1/ r - 1/ 0.043) = 0.5* 6.2*10^-3* 64^2

r = 9.7375*10^-3 m

==========

Do comment in case any doubt will reply for sure.. Goodluck